OpenStudy (anonymous):

(7^2)^3 help please ill give a medal

OpenStudy (anonymous):

117,649

OpenStudy (michele_laino):

hint: \[\large {\left( {{{\left( 7 \right)}^2}} \right)^3} = {\left( 7 \right)^{2 \times 3}}\]